chemistry

A room-temperature solution is saturated in both K
Answer #1

Answer:

1). d).Some CO2 bubbles out of solution and KCl precipitates.

2). e) 3.8 m

Molality = Molarity/density

Molality = 4/1.04

= 3.84 m

3). d).3.4*10-2

Henry's law constant (K) = C/P

= 0.12/3.5

= 0.034